Let W {(x, y, z) R3: x + y + z = 0}. Check
if W is a subspace of R3 . Find a non-zero
subspace U of R3 so that W intersection U = (0).
1
Expert's answer
2020-05-11T18:50:14-0400

a) Is W a subspace of R3?

if x + y + z = 0, that W {(-y-z, y, z) }

let's check two conditions


1) ∀ x,y∈ W, (x+y)∈ W

2) ∀ x∈W ,α∈R α*x∈W 

1)∀ a1,a2∈W

a1={x1, y1, z1}={-y1-z1, y1, z1 }

a2={x2, y2, z2}={-y2-z2, y2, z2 }

a1+a2={-y1-z1+(-y2-z2), y1+y2, z1+z2 }

x + y + z = 0:

-y1-z1+(-y2-z2)+ y1+y2+ z1+z2=0 the condition is satisfied

2)∀ a1∈W α*x∈W 

αa1={αx1, αy1, αz1}={α(-y1-z1), αy1, αz1 }

α(-y1-z1)+αy1+αz1= α((-y1-z1)+y1+z1)=0 the condition is satisfied


b) U-?

U {(x, y, z) R3: f}

where f the equation of the straight line which is passing through (0,0,0) and not lying in the x+y+z=0 plane

for example U {(x, y, z) R3: x=y=z}